When rewriting an exponential expression with a negative exponent and a positive base to an expression containing only a positve exponent does the sign of the base change? Can you answer, explain and give me an example? Explain the example so I can learn. I am realy trying hard to learn algebra in my college class.Math is my worst subject.
Answer by rfer(12730)

examples
6^-2=1/6^2=1/36
(1/4)^-2=1/(1/4)2=1/(1/16)=1/(1/16)=1*16=16
Does that help?
Bob